Sherwin Corporation has three business segments: paint, wallpapers, and tools. The company's assumed

cost of capital is 10%. Financial information about each segment follows.

Paint Segment Wallpaper Segment Tools Segment

Paint Segment

Wallpaper Segment

Tools Segment

Segment operating income

$680,000

$480,000

$800,000

Invested capital

6,000,000

3,800,000

8,000,000

(a) Calculate residual income for each segment.

(b) Assuming a reduction in assumed cost of capital to 5%, prepare a revised analysis of residual

income. Does this revised assumption alter the rankings?

A.

4(a) Paint Residual income

$90,000

4(a) Wallpaper Residual income

$110,000

4(a) Tools Residual income

$10

4(b) Paint Residual income

$390,000

4(b) Wallpaper Residual income

$295,000

4(b) Tools Residual income

$403,000

B.

4(a) Paint Residual income

$80,000

4(a) Wallpaper Residual income

$100,000

4(a) Tools Residual income

$0

4(b) Paint Residual income

$380,000

4(b) Wallpaper Residual income

$290,000

4(b) Tools Residual income

$400,000

C.

4(a) Paint Residual income

$78,000

4(a) Wallpaper Residual income

$96,000

4(a) Tools Residual income

$30

4(b) Paint Residual income

$240,000

4(b) Wallpaper Residual income

$267,000

4(b) Tools Residual income

$360,000

D.

4(a) Paint Residual income

$88,000

4(a) Wallpaper Residual income

$106,000

4(a) Tools Residual income

$40

4(b) Paint Residual income

$260,000

4(b) Wallpaper Residual income

$160,000

4(b) Tools Residual income

$240,000

E.

4(a) Paint Residual income

$66,000

4(a) Wallpaper Residual income

$76,000

4(a) Tools Residual income

$70

4(b) Paint Residual income

$385,000

4(b) Wallpaper Residual income

$293,000

4(b) Tools Residual income

$409,000

F.

None of the above
Answer #1
Residual Income = Operating income - Average Invested Assets*Required rate of return
Paint Wallpaper Tools
RI 80000 100000 0
4b 380000 290000 400000
Yes, will alter
Answer is B

B D 1 Residual Income = 0 2 Paint Wallpaper Tools 3 RI =680000-6000000*10% =480000-3800000*10% =800000 - 8000000* 10% 4 4b =6